K2[Cr(CN)2(O)2(O2) (NH3)]   oxidation state of cromium in the given complex.
 

Grade:12

1 Answers

Arun
5 years ago

K2{Cr(CN)2(O)2(O2)NH3}= k^+2 +{Cr(CN)2(O)2(O2)NH3}^-2

Now oxidation no of

CN(cyano)=-1

O(oxo)=-2

O2(peroxo)=-2

Crcromate)=x

CALCULATING OXIDATION NO OF NEGATIVE PART

x +2(-1)+2(-2)+(-2)=-2

X=2+4+2–2=+6